I need help answering part b and c of question 7

media%2F08e%2F08e2606e-0381-4043-b475-dd

I need help answering part b and c of question 7
Answer #1

b] option 4 is correct. Direction is negative x axis as opposite charges attract each other.

c] acceleration =a = F/m2 = 2.01e-5/9.5e-6   = 2.12 m/s^2